OpenStudy (anonymous):

Select the best possible first step to solving the system by first eliminating the y variable. x + 8y = –4 2x – 4y = –13 Multiply the first equation by 2. Multiply the second equation by 2. Multiply the first equation by –2. Multiply the second equation by –2.

OpenStudy (anonymous):

Multiply the second equation by 2

OpenStudy (anonymous):

Can you explain that?

OpenStudy (anonymous):

Yes multiplying second eq by 2 we will find 4x – 8y = –26 adding it in the first eq we can eliminate y x + 8y = –4 4x – 8y = –26 --------------- 5x =-30 x=-30/5 x=-6

OpenStudy (anonymous):

Thank you soooo much! :0

OpenStudy (anonymous):

You are most welcome....

OpenStudy (anonymous):

Haha! :-)